OpenStudy (mertsj):

Hey my open study isnt working and i need help. How do you convert r = 2a tan(θ)sin(θ) in polar mode to Cartesian or rectangular form? the answer is y2 = x3/(2a - x) but idk how to get that! if you help i will write you an awesome testimony! please help!!

OpenStudy (mertsj):

Posted for bridgetx516x

OpenStudy (anonymous):

\[ r=2a\tan(\theta)\sin(\theta)\\\;\\ x=r\cos(\theta)\qquad y=r\sin(\theta)\qquad\tan(\theta)={y\over x}\\ r=2a{y\over x}{y\over r}\\ r^2=2a{y^2\over x}\\ x^2+y^2={2ay^2\over x}\\ \boxed{x^3+xy^2=2ay^2} \]

OpenStudy (anonymous):

\[x^3=y^2(2a-x)\\\implies\boxed{y^2=\Large{x^3\over 2a-x}}\]

OpenStudy (anonymous):

posted for @bridgetx516x http://www-groups.dcs.st-and.ac.uk/~history/Curves/Cissoid.html

OpenStudy (mertsj):

ty on behalf of the asker

OpenStudy (anonymous):

@Mertsj I did reply in a message to the asker.

OpenStudy (mertsj):

Ty

OpenStudy (anonymous):

np